OpenStudy (anonymous):

solve: -sin^2x=2cosx-2

myininaya (myininaya):

\[-(1-\cos^2x)=2cosx-2=> -1+\cos^2x-2cosx+2=0\] \[=>\cos^2x-2cosx+1=0\]

myininaya (myininaya):

\[=>(cosx-1)^2=0\]

myininaya (myininaya):

\[cosx-1=0\]

myininaya (myininaya):

\[cosx=1\]

myininaya (myininaya):

\[x=2\pi n , n \in \mathbb{Z}\]

OpenStudy (anonymous):

thanks!